1. Find linearly independent functions that are annihilated by the given differential operator. (Give as many functions as possible. Use x as the independent variable. Enter your answers as a comma-separated list.)

D4

2. Solve the given differential equation by undetermined coefficients.
y'' + y' + y = x sin x

y(x) = ______
1
Expert's answer
2020-06-30T14:48:13-0400

Solution:

1.D4y=0

y(4)=0

"\lambda" 4=0

"\lambda" 1=0, k=4

y1(x)=1,y2(x)=x,y3(x)=x2,y4(x)=x3.

Answer: 1, x, x2, x3.


Solution:

y'' + y' + y = x sin x

y'' + y' + y = 0

"\lambda" 2+"\lambda" +1=0

"\lambda" 1=-1/2-i31/2/2, λ 2=-1/2+i31/2/2.

y0(x)=C1e-x/2cos(31/2x/2)+C2e-x/2sin(31/2x/2),

yp(x)=(Ax+B)sin(x)+(Cx+D)cos(x)

yp'(x)=(-Cx+A-D)sin(x)+(Ax+B+C)cos(x)

yp''(x)=(-Ax-B-2C)sin(x)+(-Cx+2A-D)cos(x)

xsin(x):A-C-A=1,

xcos(x):C+A-C=0,

sin(x):B+A-D-2C-B=0,

cos(x):D+B+C+2A-D=0.

=>C=-1,A=0,D=2,B=1.

yp(x)=sin(x)+(-x+2)cos(x),

y(x)=y0(x)+yp(x)

Answer:

y(x)=C1e-x/2cos(31/2x/2)+C2e-x/2sin(31/2x/2)+sin(x)+(-x+2)cos(x)